1.How much whole life insurance at age 40 can be purchased by 25 net annual premiums of RM100 each?


2.Calculate the 10th terminal reserve by the prospective method for a RM1,000 20-payment endowment insurance maturing at age 65, the policy having been issued at age 35..




1
Expert's answer
2021-01-08T12:58:53-0500

1."25\times100(60-40)=50 000"


2."20\times1000(65-35)=600 000"

"20*1000(65-45)=400 000"